OpenStudy (anonymous):

What is the value of the expression when a= 7, b= 10, c=2? 8a/b+2c

OpenStudy (eric_d):

substitute the value given

OpenStudy (kobeni-chan):

56 /10 + 4 56/14 4

OpenStudy (eric_d):

into that expression

OpenStudy (eric_d):

That's all @Lindsay_552

OpenStudy (eric_d):

Try that now

OpenStudy (ahsome):

Equation is: \[\frac{8a}{b}+2c\] Sub values: \[\frac{8*7}{10}+2*2\] \[\frac{56}{10}+4\]

OpenStudy (eric_d):

@Lindsay_552

OpenStudy (anonymous):

thanks everyone!

OpenStudy (eric_d):

\(\rlap{\color{red}{\huge\bigstar}\huge \color{blue}{ \text{You're Most Welcome! }}\color{red}\bigstar}{\; \color{aqua}{\huge\bigstar}\huge \color{aqua}{\text{You're Most Welcome! }}\color{aqua}\bigstar}\)

OpenStudy (anonymous):

so what is it